2 scientists self - fertilize a blue flowered daisy and grow 100 progeny plants that consist of 55 blue - flowered plant, 22 purple - flowered plants and 23 white - flowered plants. One of them believes this is the result of segregation of two alleles at one locus and that progeny ratio is 1:2:1. The other think the progeny phenotypes are the result of two epistatic genes that the ratio is 9:3:4. They both ask you to resolve their conflict by performing chi - square analysis on the data for both proposed mechanisms:

a. Calculate the chi - square value for hypothesis of segregation of two alleles at one locus.

b. Calculate the chi - square value for hypothesis of interaction of two epistatic genes.

c. Based on chi - square analysis, what is your conclusion regarding these two genetic hypotheses?
Answer #1

Answer 2

A) when the two allele at one locus:

Expected numbers:

Blue:(2/4) ×100=50

Purple=(1/4) ×100=25

White=(1/4) ×100=25

Chi square table:

Phenotype observed(O) Expected( E) (O-E) ^2/ E
Blue 55 50 0.5
Purple 22 25 0.36
White 23 25 0.16

Chi square experimental value: 0.5+0.36+0.16= 1.02

Now Chi square value from table at df (3-1) =2 and p=0.05 is 5.891

As Chi square experimental value is less than Chi square value from table. Our one gene with two allele hypothesis is accepted.

----

B) Intercation of two epistatic genes:

Blue= (12/16) ×100=75

Purple=(3/16) ×100=18.75

White= (1/16) ×100=6.25

Chi square table:

Phenotype Observed(O) expected (E) (O-E) ^2/E
Blue 55 75 5.33
Purple 22 18.75 0.56
White 23 6.25 44.89

Chi square experimental value is 5.33+0.56+44.89=47.87

As our Chi square experimental value is more than Chi square value from table. That's why our intercation of two epistatic gene hypothesis is rejected.

C) Based on the above result, it is found that the inheritance occurs because of one gene which contains two alleles. There is found segregation of two alleles at one locus. And the progeny ratio is found in 1:2:1.

Please rate.